# House of Cards: Massive Weights in LLMs

### v1

@@ 4.1 Zero-shot Downstream Task
Table 3: Generation tasks performance measured by Prometheus-2-7B.
Llama-3-8B LoRA 3.71 3.54 4.50 3.24 4.35 3.64 3.74 3.82
+ MacDrop 3.76 3.49 4.51 3.39 4.29 3.50 3.71 3.81
DoRA 3.85 3.71 4.56 3.34 4.35 3.59 3.73 3.80
+ MacDrop 3.79 3.41 4.55 3.26 4.29 3.79 3.85 3.85
Mistral-7B LoRA 3.55 3.30 4.61 3.33 4.41 3.45 4.05 3.81
+ MacDrop 3.75 3.42 4.58 3.26 4.35 3.30 3.98 3.81
DoRA 3.55 3.29 4.64 3.34 4.49 3.35 3.88 3.79
+ MacDrop 3.49 3.29 4.59 3.52 4.41 3.33 3.95 3.80
We fine-tune the Llama-3-8B and Mistral-7B using the alpaca_gpt4_en dataset (Peng et al., 2023) for 3 epochs (579 steps), and evaluate on five zero-shot tasks. We use two parameter-efficient fine-tuning methods, LoRA (Hu et al., 2022) and DoRA (yang Liu et al., 2024). DoRA decomposes the pre-trained weights into two components, magnitude and direction, and applies LoRA to the direction component. Our method is based on the implementation of Llama-Factory (Zheng et al., 2024) [^5] https://github.com/hiyouga/LLaMA-Factory. For MacDrop, $k$ and $p_{0}$ are set to 5 and 0.8, respectively. Details for implementations are explained in Appendix A. Table 2 presents the results on zero-shot downstream tasks. For both the models and methods, MacDrop consistently leads to performance gains, especially in ARC-Easy and ARC-Challenge tasks.
4.2 Generation Task
We evaluate on the generated texts of the same models in Section 4.1 using the Spec-Bench dataset (Xia et al., 2024). This benchmark includes six subtasks, each containing 80 instances: multi-turn (MT) conversation from MT-bench (Zheng et al., 2023), translation from WMT14 DE-EN (Bojar et al., 2014), summarization from CNN/Daily Mail (Nallapati et al., 2016), question answering (QA) from Natural Questions (Kwiatkowski et al., 2019), mathematical reasoning from GSM8K (Cobbe et al., 2021), and retrieval-augmented generation (RAG) from Natural Questions (Kwiatkowski et al., 2019). We utilize the direct assessment of Prometheus-2-7B (Kim et al., 2024) to evaluate the generated texts using a 5-point Likert scale. Prometheus-2-7B is an open-source language model specifically designed for evaluation purposes [^6] https://github.com/prometheus-eval/prometheus-eval. Table 3 presents the results on generation tasks. MT-1 and MT-2 indicate the first turn and second turn, respectively. Unfortunately, MacDrop shows limited performance improvements in generation tasks. Examples of the generated texts and judgements are provided in Appendix E.
@@ 4.3 Ablation Study
We further provide ablation studies related to MacDrop using Llama-3-8B. Unless otherwise stated, for MacDrop, $k$ and $p_{0}$ are set to 5 and 0.8.
@@ 4.3.1 Dropout Scope and Probability
Figure 3(b) illustrates the mean zero-shot accuracy according to the dropout scope and dropout probability $p$. It is observed that among three scopes, the original performance (i.e., without dropout), represented by the dotted line at 76.0, can be surpassed only when dropout is applied solely to massive weights. Nevertheless, if strong dropout (e.g., $p\geq 0.85$) is maintained on the pre-trained massive weights during fine-tuning, performance deteriorates. This highlights the need to safeguard the pre-trained massive weights during the final stages of fine-tuning, because we are ultimately using them without causing any damage.
@@ 4.3.2 Curriculum Methods and Initial Dropout Probability
5 Related work
Nevertheless, there has been limited in-depth research on the phenomenon itself. In fact, the idea of global attentions, such as [CLS] and [SEP] tokens–similar to attention sinks–was introduced and emphasized even before the LLM era (Zaheer et al., 2020; Beltagy et al., 2020). In the LLM era, Yu et al. (2024) showed that sink tokens can appear not only at the beginning of a sentence but also in the middle, and they are often shown to be nonsemantic (e.g., ‘.’). Sun et al. (2024) discovered the presence of massive activations in the hidden state space of sink tokens, demonstrating that massive activations trigger the attention sinks phenomenon. Meanwhile, in vision transformers, similar phenomenon is observed (Darcet et al., 2024). They showed that training with register tokens, which is additional meaningless tokens similar to sink tokens, resulted in improved dense prediction and interpretability. Different from previous work, we explore this phenomenon in the weight space.
@@ 6 Conclusion
In this paper, we explore the weight space of LLMs and identify the presence of massive weights within a FFN module in an early layer, which are predominantly pre-trained and have a significant impact on the performance of LLMs. Based on our observation, we propose a plug-and-play fine-tuning method called MacDrop, which applies dropout to the pre-trained massive weights, rather than to the parameters of adapters, during parameter-efficient fine-tuning. We hope that our findings will inspire future research in weight space of LLMs, including model merging (Li et al., 2023) and model editing (Yao et al., 2023).
Reproducibility
For our analysis in Section 2, we conduct a consistent and reproducible analysis using only the bos token (Section 2.2) and provide the specific position of massive weights across various models in Appendix B. For our algorithm, MacDrop, we provide PyTorch-style pseudo code in Section 3 and training details in Appendix A. Furthermore, we present github links for all our implementations with footnotes, when necessary.
@@ Appendix B Position of Massive Weights
Table 6 summarizes the position of massive weights across various models.
@@
gemma-2-2b - -
gemma-2-2b-it - -
gemma-2-9b - -
gemma-2-9b-it - -
gemma-2-27b - -
gemma-2-27b-it - -
Appendix C Futher Analysis for Various LLMs
@@ C.2 Llama-3 family
C.3 Mistral and Mixtral family
@@ C.5 Gemma-2 family
Appendix D Attention Sinks
Figure 13 describes attention after Softmax in the early layers (from layer 1 to layer 8) across various models. Attention sinks are observed in the layers after the massive weights layer. In Llama-2-7B (Figure 13(a)), Mistral-7B (Figure 13(c)), and Mixtral-8x7B (Figure 13(d)), sink tokens are the initial token (‘Summer’) and the first delimiter token (‘.’), discovered by Sun et al. (2024). In Llama-3-8B (Figure 13(b)) and Phi-3-mini (Figure 13(e)), sink token is the only the initial token (‘Summer’). Interestingly, in these five models, it is commonly observed that significant attention is concentrated on non-semantic tokens (‘.’) before attention sinks occur. However, in Gemma-2 (Figure 13(f)), attention sinks do not happen and attention is primarily assigned to local tokens. Note that what we provide is the average of the heads, and there might be heads that do not fully sink when viewed individually.
@@ Appendix D Attention Sinks
Appendix E Examples in Generation Task
Assistant A and B indicate Llama-3-8B with and without MacDrop, respectively.

### v2

@@ 3 Massive Weights Curriculum Dropout
4 Experiments
4.1 Zero-shot Downstream Task
We fine-tune the Llama-3-8B and Mistral-7B using the alpaca_gpt4_en dataset (Peng et al., 2023) for 3 epochs (579 steps), and evaluate on five zero-shot tasks. We use two parameter-efficient fine-tuning (PEFT) methods, LoRA (Hu et al., 2022) and DoRA (yang Liu et al., 2024). DoRA decomposes the pre-trained weights into two components, magnitude and direction, and applies LoRA to the direction component. Our method is based on the implementation of Llama-Factory (Zheng et al., 2024) [^5] https://github.com/hiyouga/LLaMA-Factory. For MacDrop, $k$ and $p_{0}$ are set to 5 and 0.8, respectively. Details for implementations are explained in Appendix A.
Table 3 presents the results on zero-shot downstream tasks. For both the models and methods, MacDrop consistently leads to performance gains, especially in ARC-Easy and ARC-Challenge tasks. Note that we do not claim MacDrop is superior to LoRA/DoRA; rather, we demonstrate that LoRA/DoRA performs better when MacDrop is applied than when it is not. Results for other LLMs are provided in Appendix F.
MacDrop is designed to mitigate dependence on massive weights during PEFT. To assess whether this dependency is effectively reduced, we attack the previous fine-tuned models in Table 3. Table 4 shows the performance changes whether applying MacDrop under the top-3 zeroing attack. The zeroing attack severely degrades the performance of fine-tuned models without MacDrop, as seen in the case of Llama-3-8B with DoRA, where performance drops from 76.1 to 40.3. However, fine-tuned models with MacDrop exhibit significantly better performance under attack, indicating better robustness. Especially, when MacDrop is combined with DoRA, it demonstrates remarkable robustness. For instance, in the case of Llama-3-8B with DoRA, where performance drops from 76.4 to 72.7.
Table 4: Zero-shot downstream tasks performance under the top-3 zeroing attack, related to robustness.
Model Method ARC-E ARC-C BoolQ PIQA WG Avg.
Llama-3-8B LoRA 29.9 22.9 45.7 52.9 50.7 40.4
+ MacDrop 36.8 24.7 64.3 58.5 54.1 47.7
DoRA 29.8 23.0 46.0 52.6 50.0 40.3
+ MacDrop 78.7 53.7 79.9 79.0 72.4 72.7
Mistral-7B LoRA 54.6 34.8 58.0 74.8 57.5 55.9
+ MacDrop 69.2 45.0 78.3 79.3 64.1 67.2
DoRA 55.7 35.3 58.6 74.4 58.2 56.4
+ MacDrop 78.1 52.4 84.1 82.3 69.5 73.3
4.2 Long Context Task
We evaluate the two LLMs on LongBench (Bai et al., 2024), a benchmark specifically designed to assess the ability to understand long contexts. This includes 5 sub-categories and 16 English datasets: single-document QA, multi-document QA, summarization, few-shot learning, synthetic, and code generation. We set the max length of models to 7,500. Table 2 shows that MacDrop increases the performance when understanding long context is required.
@@ 4.3 Ablation Study
We further provide ablation studies related to MacDrop using Llama-3-8B. The variations used in this section can serve as a baseline.
@@ 4.3.1 Dropout Scope and Probability
Figure 6 illustrates the mean zero-shot accuracy according to the dropout scope and dropout probability $p$. It is observed that among three scopes, the original performance (i.e., without dropout), represented by the dotted line at 76.0, can be surpassed only when dropout is applied solely to massive weights. Nevertheless, if strong dropout (e.g., $p\geq 0.85$) is maintained on the pre-trained massive weights during fine-tuning, performance deteriorates. This highlights the need to safeguard the pre-trained massive weights during the final stages of fine-tuning, because we are ultimately using them without causing any damage.
@@ 5 Conclusion
In this paper, we explore the weight space of LLMs and identify the presence of massive weights within a FFN module in an early layer, which are predominantly pre-trained and have a significant impact on the performance of LLMs. Based on our observation, we propose a plug-and-play fine-tuning method called MacDrop, which applies dropout to the pre-trained massive weights, rather than to the parameters of adapters, during parameter-efficient fine-tuning. We hope that our findings will inspire future research in weight space of LLMs.
Impact Statement
This paper seeks to contribute to the advancement of large language models, particularly by providing deeper insights into their internal mechanisms. Our research demonstrates, through various open-sourced LLMs, that in many cases, models can be easily disrupted by zeroing massive weights if access to their weights is available. Subsequently, we improve robustness against such attacks through MacDrop. The existence of massive weights does not inherently present a risk if the model weights remain undisclosed. However, we believe that model developers should understand these internal mechanisms and take responsibility for ensuring robustness. In this vein, our work has the potential to make a meaningful impact.
@@ Appendix B Position of Massive Weights
Table 7 summarizes the position of massive weights across various models. These are selected based on the magnitudes of intermediate state in Appendix D.
@@
gemma-2-2b 2 [1257, 2896, 6954, 8624, 7118]
gemma-2-2b-it 2 [1257, 2896, 6954, 8624, 9140]
gemma-2-9b 1 [2769, 6656, 4889, 14293, 11065]
gemma-2-9b-it 1 [2769, 6656, 4889, 14293, 10429]
gemma-2-27b 10 [34659, 32862, 9590, 8959, 32744]
gemma-2-27b-it 10 [34659, 32862, 9590, 32744, 8959]
Appendix C bos Token Analysis for Various LLMs
In this section, we provide the magnitudes of activations of the hidden state and normalized attention scores according to the position of the bos token, after massive activations appear (specifically, in the middle layer), for various LLM families.
C.1 Llama-2 family
Llama-2-7B (Figure 8) has massive activations at the starting token or first delimiter token (first column). When the bos token is placed in the starting position, it triggers massive activations and the ‘Summer’ token loses its massive activations, while first delimiter token ‘.’ still keeps its massive activations (second column). When the bos token is placed in the middle or ending position after the first delimiter token, it does not trigger massive activations (third and fourth columns).
Llama-2-13B (Figure 9) has massive activations only at the starting token, other than Llama-2-7B (first column). In cases where the bos token is inserted, the same tendencies are observed as with the LLaMA-2-7B model.
Figure 8: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Llama-2-7b-hf.
Figure 9: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Llama-2-13b-hf.
C.2 Llama-3 family
Llama-3-8B (Figure 10) does not have massive activations at delimiter tokens such as ‘.’ (first column). When the bos token is placed in the starting position, it triggers massive activations and the ‘Summer’ token loses its massive activations, similar to Llama-2 family (second column). When the bos token is placed in the middle or ending position, it also triggers massive activations in the same feature dimensions (third and fourth columns). Namely, the bos token has massive activations, regardless of its position. What is intriguing is that, despite the difference in magnitude according to the position, the bos token similarly exhibits attention sinks.
Llama-3-70B (Figure 11) generally exhibits similar trends to Llama-3-8B. One notable difference is that the degree of sinking for the token at the first position is significantly stronger compared to that of the Llama-3-8B.
Figure 10: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Meta-Llama-3-8B.
Figure 11: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Meta-Llama-3-70B.
C.3 Mistral and Mixtral family
Mistral (Figure 12) does not exhibit massive activations at the starting position and does not trigger attention sink, contrary to previous findings observed by Sun et al. (2024a). Rather, massive activations are observed only at the first delimiter token (first column). When the bos token is placed in the starting position, the first delimiter token loses its massive activations (second column). However, when the bos token is placed in the middle or ending position after the first delimiter token, massive activations are observed in both tokens (third and fourth columns). Similar to Llama-3 family, the bos token has massive activations, regardless of its position.
Mixtral (Figure 13) exhibits the same behavior as Mistral. The only difference is observed in the magnitude of its massive activations, with Mixtral producing values approximately ten times higher than Mistral.
Figure 12: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Mistral-7B-v0.1.
Figure 13: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Mixtral-8x7B-v0.1.
C.4 Phi-3 family
Phi-3-mini (Figure 14) and Phi-3-medium (Figure 15) exhibit a similar tendency to Llama-2-13B. This family has massive activations only at the starting token (first column), with a similar response when the bos token is inserted (second, third, and fourth column). A significant distinction between the Llama-2-13B model and the Phi-3 family lies in their attention mechanisms. Specifically, the Phi-3 family demonstrates weaker attention on the token at the first position than Llama-2-13B model. This reduced attention appears to be primarily redistributed to recent tokens.
Figure 14: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Phi-3-mini-4k-instruct.
Figure 15: (Top) Magnitudes of the hidden state and (Bottom) attention scores of Phi-3-medium-4k-instruct.
C.5 Gemma-2 family
The Gemma-2 family displays significantly distinct magnitudes of activations and attention scores when compared to other model families. This divergence remains consistent regardless of the absence (first column) or presence (other columns) of the bos token.
Gemma-2-2b (Figure 16) and Gemma-2-9b (Figure 17) do not exhibit noticeably large values along either the token axis or the feature dimension axis, from the perspective of magnitudes of activations (first column, top). This suggests that massive activations are not present. As a result, the attention mechanism avoids the attention sink phenomenon and demonstrates a strong attention on the locality of recent tokens (first column, bottom). However, when the bos token is fed into these models, it exhibits massive activations with extremely large values in certain feature dimensions, regardless of its position (second, third, and fourth columns). Nevertheless, compared to other models where attention sinks occur, they allocate significantly greater attention to recent tokens (especially, to its own tokens).
Figure 16: (Top) Magnitudes of the hidden state and (Bottom) attention scores of gemma-2-2b.
Figure 17: (Top) Magnitudes of the hidden state and (Bottom) attention scores of gemma-2-9b.
Gemma-2-27b (Figure 18) demonstrates a distinct behavior compared to smaller models. It exhibits noticeably large values along the feature dimension axis across all tokens, from the perspective of magnitudes of activations (first column, top). This distribution, where the differences between tokens are not pronounced, fails to create attention sinks (first column, bottom). When the bos token is placed in the starting position, it triggers massive activations and attention sinks by generating value that exceed the magnitudes of other tokens by more than tenfold, in the certain feature dimension (second column). However, when the bos token is placed in the middle or ending position, it does not trigger massive activations, similar to when the bos token is absent (third and fourth columns).
Figure 18: (Top) Magnitudes of the hidden state and (Bottom) attention scores of gemma-2-27b.
In summary,
  • Llama-2, Llama-3, and Phi-3 families have massive activations at the first position.
  • Llama-3, Mistral/Mixtral families, and Gemma-2-2/9b models have massive activations at the bos token.
  • All families have massive activations at the bos token placed at the first position.
Appendix D Futher Analysis for Various LLMs
@@ D.2 Llama-3 family
D.3 Mistral and Mixtral family
@@ D.5 Gemma-2 family
Appendix E Attention Sinks
Figure 24 describes attention after Softmax in the early layers (from layer 1 to layer 8) across various models. Attention sinks are observed in the layers after the massive weights layer. In Llama-2-7B (Figure 24(a)), Mistral-7B (Figure 24(c)), and Mixtral-8x7B (Figure 24(d)), sink tokens are the initial token (‘Summer’) and the first delimiter token (‘.’), discovered by Sun et al. (2024a). In Llama-3-8B (Figure 24(b)) and Phi-3-mini (Figure 24(e)), sink token is the only the initial token (‘Summer’). Interestingly, in these five models, it is commonly observed that significant attention is concentrated on non-semantic tokens (‘.’) before attention sinks occur. However, in Gemma-2 (Figure 24(f)), attention sinks do not happen and attention is primarily assigned to local tokens. Note that what we provide is the average of the heads, and there might be heads that do not fully sink when viewed individually.
@@ Appendix E Attention Sinks
Appendix F Zero-shot Downstream Task across Various LLMs
We address the potential limitations of algorithms based on massive phenomena. Table 8 presents the results on zero-shot downstream tasks across different LLMs. The results show that MacDrop is not effective for LLMs that are not sensitive to massive weights, such as Phi-3-medium and Gemma-2 family, as shown in Section 2.
Nevertheless, we believe that clearly demonstrating such limitations is itself meaningful. Most previous and ongoing research addressing massive phenomena have stated that various models exhibit massive phenomena, and have proposed algorithms and experimental results excluding Gemma-2 family. Because we conduct experiments on a wide variety of recent LLMs, we are able to reveal such limitations. Additionally, we would like to emphasize that, similar to previous studies, MacDrop demonstrates sufficient feasibility for LLMs exhibiting massive phenomena.
Table 8: Zero-shot downstream tasks performance across different LLMs.
Model Method ARC-Easy ARC-Challenge BoolQ PIQA WinoGrande Avg.
Llama-2-13B LoRA 77.1 51.9 82.2 81.6 72.8 73.1
+ MacDrop 78.8 52.6 82.2 71.4 71.9 73.4
DoRA 77.4 51.8 82.0 81.6 72.8 73.1
+ MacDrop 77.8 52.5 81.9 81.8 72.3 73.3
Llama-3-8B LoRA 79.6 58.2 83.9 82.4 75.9 76.0
+ MacDrop 82.9 58.3 83.9 82.6 75.0 76.5
DoRA 80.8 57.7 83.9 82.5 75.8 76.1
+ MacDrop 81.9 58.2 83.9 82.2 75.6 76.4
Llama-3-70B LoRA 87.5 66.3 86.1 84.9 80.4 81.0
+ MacDrop 87.5 66.5 86.1 85.0 80.8 81.2
DoRA 87.5 66.6 86.1 85.0 80.8 81.2
+ MacDrop 87.4 66.6 86.0 85.1 80.9 81.2
Mistral-7B LoRA 78.5 54.9 84.9 82.9 75.3 75.3
+ MacDrop 80.9 56.7 85.0 83.0 75.3 76.2
DoRA 78.4 55.1 85.0 82.9 75.1 75.3
+ MacDrop 80.6 56.7 85.3 82.9 75.1 76.1
Phi-3-mini LoRA 72.5 53.7 86.4 80.1 74.0 73.3
+ MacDrop 75.0 54.7 86.2 80.5 74.0 74.1
DoRA 72.3 53.3 86.4 80.0 73.6 73.1
+ MacDrop 72.9 53.5 86.3 80.0 73.7 73.3
Phi-3-medium LoRA 81.4 62.2 88.7 82.6 76.4 78.3
+ MacDrop 81.2 61.9 88.7 82.5 76.4 78.1
DoRA 80.92 62.0 88.6 82.4 76.2 78.0
+ MacDrop 80.8 61.9 88.5 82.4 76.2 77.9
Gemma-2-2b LoRA 81.6 54.4 79.8 79.2 68.7 72.7
+ MacDrop 81.5 54.2 79.6 79.2 68.6 72.6
DoRA 81.4 54.0 79.6 79.3 68.7 72.6
+ MacDrop 81.6 53.9 79.4 79.3 68.7 72.6
Gemma-2-9b LoRA 89.6 69.0 86.5 82.8 75.3 80.6
+ MacDrop 89.2 68.4 86.5 82.8 75.2 80.4
DoRA 89.4 68.9 86.2 82.7 75.8 80.6
+ MacDrop 89.2 68.5 86.3 82.8 75.7 80.5
Gemma-2-27b LoRA 87.5 69.0 86.0 84.1 80.6 81.4
+ MacDrop 87.3 68.8 86.0 84.2 80.5 81.4
DoRA 88.5 69.4 85.6 84.5 80.0 81.6
+ MacDrop 88.1 69.2 85.7 84.3 80.0 81.5
Appendix G Generation Tasks
We evaluate on the generated texts of the same models in Section 4.1 using the Spec-Bench dataset (Xia et al., 2024). This benchmark includes six subtasks, each containing 80 instances: multi-turn (MT) conversation from MT-bench (Zheng et al., 2023), translation from WMT14 DE-EN (Bojar et al., 2014), summarization from CNN/Daily Mail (Nallapati et al., 2016), question answering (QA) from Natural Questions (Kwiatkowski et al., 2019), mathematical reasoning from GSM8K (Cobbe et al., 2021), and retrieval-augmented generation (RAG) from Natural Questions (Kwiatkowski et al., 2019). We utilize the direct assessment of Prometheus-2-7B (Kim et al., 2024) to evaluate the generated texts using a 5-point Likert scale. Prometheus-2-7B is an open-source language model specifically designed for evaluation purposes [^6] https://github.com/prometheus-eval/prometheus-eval. Table 9 presents the results on generation tasks. MT-1 and MT-2 indicate the first turn and second turn, respectively. Unfortunately, MacDrop shows limited performance improvements in generation tasks. Table 10 provides examples of the generated texts and judgements. Assistant A and B indicate Llama-3-8B with and without MacDrop, respectively.
Table 9: Generation tasks performance measured by Prometheus-2-7B.
Llama-3-8B LoRA 3.71 3.54 4.50 3.24 4.35 3.64 3.74 3.82
+ MacDrop 3.76 3.49 4.51 3.39 4.29 3.50 3.71 3.81
DoRA 3.85 3.71 4.56 3.34 4.35 3.59 3.73 3.80
+ MacDrop 3.79 3.41 4.55 3.26 4.29 3.79 3.85 3.85
Mistral-7B LoRA 3.55 3.30 4.61 3.33 4.41 3.45 4.05 3.81
+ MacDrop 3.75 3.42 4.58 3.26 4.35 3.30 3.98 3.81
DoRA 3.55 3.29 4.64 3.34 4.49 3.35 3.88 3.79
+ MacDrop 3.49 3.29 4.59 3.52 4.41 3.33 3.95 3.80
@@ Appendix G Generation Tasks
Appendix H Related work
Nevertheless, there has been limited in-depth research on the phenomenon itself. In fact, the idea of global attentions, such as [CLS] and [SEP] tokens–similar to attention sinks–was introduced and emphasized even before the LLM era (Zaheer et al., 2020; Beltagy et al., 2020). In the LLM era, Yu et al. (2024b) showed that sink tokens can appear not only at the beginning of a sentence but also in the middle, and they are often shown to be nonsemantic (e.g., ‘.’). Sun et al. (2024a) discovered the presence of massive activations in the hidden state space of sink tokens, demonstrating that massive activations trigger the attention sinks phenomenon. Meanwhile, in vision transformers, similar phenomenon is observed (Darcet et al., 2024). They showed that training with register tokens, which is additional meaningless tokens similar to sink tokens, resulted in improved dense prediction and interpretability. Different from previous work, we explore this phenomenon in the weight space.
Specifically, we define the massive weights in an activation-aware manner using the bos token. Similarly, Wanda (Sun et al., 2024b) with a structured pruning (An et al., 2024) and AWQ (Lin et al., 2024) calculate weight importance scores based on a small of calibration data. However, it is important to note that the massive weights are confined to a specific single layer, whereas Wanda and AWQ identify important weights within every linear layer. In other words, the massive weights would be included among those selected through Wanda or AWQ. Our contribution focuses more deeply on a narrowly defined aspect compared to these studies.
Concurrently with our work, Yu et al. (2024a) have defined “super weight,” which is a single weight that can significantly influence the performance of LLMs. To compare more specifically, we identify that the intermediate state in the layer $l$ possesses massive activations. Accordingly, we define massive weight as the corresponding rows of ${\bm{W}}_{up}$. However, Yu et al. (2024a) take into account both the intermediate state and the subsequent hidden state. In other words, they define super weight as the intersection of the corresponding column of ${\bm{W}}_{down}$ to the super activation in the intermediate state and the corresponding row of ${\bm{W}}_{down}$ to the super activation in the hidden state, at the layer $l$. In summary, we examine the intermediate state exclusively from an output perspective (i.e., ${\bm{W}}_{up}$), while their approach appears to have considered it from an input perspective (i.e., ${\bm{W}}_{down}$), incorporating the hidden state.